ok i have math hw and one of the problems says this X+(-9)+9=8
the next problem says X+(-2.8)+9.2=0 i dont understand how to do these kinds of problems and we don't have the right textbook ?????

Answer by Alan3354(69443)

X+(-9)+9=8
+9 and -9 add to zero
x + 0 = 8
x = 8
----------
the next problem says X+(-2.8)+9.2=0
9.2 - 2.8 = 6.4
x + 6.4 = 0
Subtract 6.4
x = -6.4